# Average Calibration Losses for Reliable Uncertainty in Medical Image Segmentation

### v2

@@ 6.3 Presence of Missing Ground Truth Classes
Another critical observation is the significant impact that images with missing ground truth classes can have on calibration performance. During both training and evaluation, the inclusion of cases with missing classes tends to skew image-level calibration error metrics, leading to an apparent decrease in calibration improvement. By excluding missing classes from the calibration error computation during training, as we have presented here, a more consistent improvement in calibration performance is observed. Similarly, by limiting evaluation to cases with full ground truth classes we observe a more consistent improvement in calibration metrics, however, for completeness we primarily report results on the entire testing dataset. The softening of probability distributions towards the decision boundary, an effect induced by our auxiliary losses, can inadvertently penalise predictions when a class is entirely absent. This observation represents a limitation of our method and suggests that the utility of these auxiliary losses may be reduced in scenarios where missing classes are prevalent. Although this issue is not specific for our calibration losses, as it is also shown to affect Dice loss [41].
@@ 6.4 Comparison of Hard and Soft Binning
When comparing the hard- and soft-binned calibration losses, our results indicate that the soft-binned variant generally delivers superior calibration performance. However, this improvement in calibration comes at a cost: soft binning results in a statistically significant reduction in segmentation performance, as measured by the Dice score, across all datasets except KiTS. This trade-off between calibration and segmentation performance is most pronounced with the soft-binned approach. Conversely, the hard-binned calibration loss maintains segmentation performance comparable to the baseline, with the exception of a modest 0.2% decrease in Dice score for the AMOS 22 dataset, suggesting that it may be more appropriate in applications where segmentation accuracy is paramount.
@@ 6.5 Limitations of Calibration Losses
Although our auxiliary calibration losses lead to marked improvements in calibration metrics, it remains to be demonstrated that these gains translate to better clinical outcomes. It is perhaps not surprising that losses based on established calibration metrics would improve those metrics during evaluation. However, our approach is notable for allowing these metrics to be computed at an image level and directly incorporating these as auxiliary losses in a semantic segmentation context. This strategy leverages the large capacity of deep neural networks to learn better-calibrated predictions while exerting only limited negative effects on segmentation performance. It is important to emphasise that our comparisons are made against a baseline loss comprising both DSC and cross-entropy (CE), with the latter generally producing better-calibrated results. In contrast, our previous work relied on a DSC loss as a comparator, which yields highly miscalibrated predictions. One inherent limitation of our current method is that the process of binning foreground probabilities strips away spatial information, thereby limiting the model’s ability to capture spatial correlations in uncertainty and confidence. Although the visualisations in Fig. 5 demonstrate instances where improvements in calibration align with error regions, this correspondence is not consistent across all cases.
@@ 6.6 Conclusion
In this study, we introduced differentiable formulations of marginal L1 Average Calibration Error (mL1-ACE) as auxiliary losses specifically designed for improving the calibration of deep neural networks in medical image segmentation tasks. Through comprehensive evaluations on four publicly available datasets (ACDC, AMOS, BraTS, KiTS), we demonstrated that integrating calibration-aware losses at training time significantly reduces calibration errors, as evidenced by substantial improvements in both macro- and micro-averaged ACE and MCE metrics. The soft-binned mL1-ACE variant provided the largest calibration improvements, reducing macro- and micro-averaged ACE by, an average of, 32% and 48%, respectively, at a small cost in segmentation accuracy (-0.9% DSC). In contrast, the hard-binned variant maintained segmentation accuracy (+0.06% DSC) while reducing macro- and micro-averaged ACE by 16% and 35%, respectively.

### v3

@@ 6.3 Presence of Missing Ground Truth Classes
Another critical observation is the significant impact that images with missing ground truth classes can have on calibration performance. During both training and evaluation, the inclusion of cases with missing classes tends to skew image-level calibration error metrics, leading to an apparent decrease in calibration improvement. By excluding missing classes from the calibration error computation during training, as we have presented here, a more consistent improvement in calibration performance is observed. Similarly, by limiting evaluation to cases with full ground truth classes we observe a more consistent improvement in calibration metrics, however, for completeness we primarily report results on the entire testing dataset. The softening of probability distributions towards the decision boundary, an effect induced by our auxiliary losses, can inadvertently penalise predictions when a class is entirely absent. This observation represents a limitation of our method and suggests that the utility of these auxiliary losses may be reduced in scenarios where missing classes are prevalent. Although this issue is not specific for our calibration losses, as it is also shown to affect DSC loss [44].
@@ 6.4 Comparison of Hard and Soft Binning
When comparing the hard- and soft-binned calibration losses, our results indicate that the soft-binned variant generally delivers superior calibration performance, but at the cost of a statistically significant reduction in segmentation performance (DSC) on all datasets except KiTS.
Quantitatively, hard-binning (hL1-ACE) provides moderate calibration improvements (average 16% ACE reduction) while essentially maintaining segmentation performance (average +0.03 percentage points DSC). In contrast, soft-binning (sL1-ACE) achieves larger calibration gains (average 33% ACE reduction) but with a small decrease in segmentation accuracy (average -0.7 percentage points DSC), highlighting a systematic calibration–accuracy trade-off.
Conversely, the hard-binned calibration loss maintains segmentation performance comparable to the baseline, with only a modest DSC decrease on AMOS 22, suggesting that it may be more appropriate in applications where segmentation accuracy is paramount. The choice between variants should depend on application requirements: hard-binning is preferable when segmentation accuracy is paramount and some calibration improvement is desired; soft-binning is preferable when calibration is critical (e.g., for safety-critical applications, active learning, or quality control) and the small segmentation trade-off is acceptable.
The sensitivity analysis further clarifies these trade-offs. Hard L1-ACE is comparatively robust to both bin count and loss weighting: DSC remains stable across a wide range of settings, while calibration steadily improves as the ACE weighting increases, making it a suitable choice when segmentation accuracy is the primary objective and hyperparameter tuning capacity is limited. Soft L1-ACE, in contrast, achieves stronger calibration improvements but is more sensitive to hyperparameters: larger numbers of bins and higher ACE weightings lead to noticeable reductions in DSC, reflecting a sharper accuracy–calibration trade-off and suggesting the use of coarser binning and moderate or lower ACE weightings (e.g., $\lambda<1.0$) when segmentation quality must be preserved.
The dynamics of joint optimisation explain why calibration and segmentation performance are inherently in tension. Overlap metrics such as DSC reward sharp, confident predictions at correct spatial locations, whereas calibration losses are spatially agnostic and depend only on the alignment between predicted confidence and empirical accuracy. The soft-binned variant achieves superior calibration because its overlapping bin assignments (Fig. 2) impose stronger smoothing on class probabilities, more effectively reducing overconfidence but also blurring decision boundaries and thereby degrading DSC. The hard-binned variant, with its discrete binning, preserves sharper probability distributions and thus segmentation accuracy, while still improving calibration through explicit optimisation of the ACE objective. In practice, this means that Hard L1-ACE is a robust default for general use, whereas Soft L1-ACE is best deployed when improved calibration justifies more careful hyperparameter tuning and a modest reduction in DSC.
6.5 Comparison with Temperature Scaling and Cross-Entropy Only
Post-hoc temperature scaling provides a simple and effective way to improve calibration for the DSC+CE baseline, and our results confirm meaningful gains using this approach. In comparison, Hard mL1-ACE achieves calibration and DSC performance that is broadly comparable to temperature scaling, while Soft mL1-ACE provides superior calibration across datasets. Local Temperature Scaling [13] may offer further post-hoc improvements for dense prediction tasks, but such methods still require a separate calibration phase, whereas our mL1-ACE losses integrate calibration directly into the training objective.
Cross-entropy alone produces better-calibrated probabilities than DSC+CE but does so at the expense of segmentation accuracy, illustrating the well-known trade-off between these objectives [29, 30]. Both hard and soft mL1-ACE variants achieve superior calibration compared to CE-only while preserving the strong segmentation performance provided by DSC. This demonstrates that our auxiliary losses mitigate the calibration–accuracy trade-off by enabling simultaneous improvement of both properties, an important consideration for reliable clinical deployment.
6.6 Practical Implications
Our work demonstrates that the use of our auxiliary losses makes the class-wise probabilities more reliable as indicators of accuracy. While we cannot definitively quantify the extent to which a given reduction in calibration error translates to improved clinical outcomes without further downstream studies, the improved reliability of confidence estimates has several potential practical applications. Models trained with auxiliary calibration losses could be integrated into quality control systems to automatically flag uncertain predictions that are likely to be inaccurate, thereby streamlining the review process for radiologists. Furthermore, reliable uncertainty estimates are crucial for active learning scenarios, where they can guide sample selection to maximise data efficiency. Finally, providing clinicians with trustworthy confidence information alongside segmentations can enhance trust in AI systems and support more informed decision-making. However, further studies are needed to verify these impacts in clinical settings.
@@ 6.7 Limitations of Calibration Losses
Although our auxiliary calibration losses lead to marked improvements in calibration metrics, it remains to be demonstrated that these gains translate to better clinical outcomes. It is perhaps not surprising that losses based on established calibration metrics would improve those metrics during evaluation. However, our approach is notable for allowing these metrics to be computed at an image level and directly incorporating these as auxiliary losses in a semantic segmentation context. This strategy leverages the large capacity of deep neural networks to learn better-calibrated predictions while exerting only limited negative effects on segmentation performance. One inherent limitation of our current method is that the process of binning foreground probabilities strips away spatial information, thereby limiting the model’s ability to capture spatial correlations in uncertainty and confidence. Although the visualisations in Fig. 5 demonstrate instances where improvements in calibration align with error regions, this correspondence is not consistent across all cases.
A further avenue for future work concerns the binning kernels themselves. In this study, we employed fixed square and triangular kernels for hard- and soft-binning, respectively. Learning the kernel shape during training, for example, by parameterising the bin membership functions and optimising them jointly with the network weights, similar to [22], could allow the model to adapt to dataset-specific probability distributions.
Recent work has highlighted important considerations for calibration in deep learning. Chidambaram & Ge discuss the calibration-generalization trade-off and methods for fair comparison of calibration approaches [10]. Wang et al. argue that some degree of overconfidence may be acceptable or even beneficial in certain contexts [45]. Our work contributes to this discussion by providing practitioners with explicit control over the calibration objective, enabling them to navigate the accuracy-calibration trade-off based on their application’s specific requirements.
Our evaluation across four diverse datasets, covering different modalities (MRI, CT) and anatomical regions, provides strong evidence of our method’s reliability. Dedicated cross-dataset generalisation experiments, where models trained on one dataset are evaluated on another, would be a valuable direction for future work to further assess robustness under severe distribution shifts.
@@ 6.8 Conclusion
In this study, we introduced differentiable formulations of marginal L1 Average Calibration Error (mL1-ACE) as auxiliary losses specifically designed for improving the calibration of deep neural networks in medical image segmentation tasks. Through comprehensive evaluations on four publicly available datasets (ACDC, AMOS, BraTS, KiTS), we demonstrated that integrating calibration-aware losses at training time significantly reduces calibration errors, as evidenced by substantial improvements in both macro- and micro-averaged ACE and MCE metrics. The soft-binned mL1-ACE variant provided the largest calibration improvements, reducing macro- and micro-averaged ACE by an average of 33% and 51%, respectively, at a small cost in segmentation accuracy (-0.7% DSC). In contrast, the hard-binned variant maintained segmentation accuracy (+0.03% DSC) while reducing macro- and micro-averaged ACE by 16% and 32%, respectively.
Importantly, our comprehensive baseline comparisons demonstrate that both mL1-ACE variants outperform not only the standard DSC+CE baseline but also CE-only training in terms of calibration, across almost all metrics. While CE alone produces better-calibrated predictions than DSC+CE, this comes at a substantial cost to segmentation performance. Our approach successfully resolves this fundamental trade-off, enabling practitioners to achieve both excellent segmentation accuracy and superior calibration without forcing a choice between these objectives. This represents a significant practical advance for medical imaging applications.